\begin{thebibliography}{99}
\bibitem{AB1}  M. Abbas and B. E. Rhoades, \emph{Fixed and periodic point results in cone metric space}, Appl. Math. Lett., \textbf{22}(4) (2009), 511--515.

\bibitem{AB2} M. Abbas, G. Jungck, \emph{Common ﬁxed point results for noncommuting mappings without continuity in cone metric spaces}, J. Math. Anal. Appl., \textbf{341} (2008), 416--420.

\bibitem{AY} H. Aydi, M. Abbas and C. Vetro, \emph{Partial Hausdorff metric and Nadler's fixed point theorem on partial metric spaces}, Top. Appl., 159 (2012), 3234--3242.

\bibitem{BA}  S. Banach, \emph{Sur les operations dans les ensembles abstraits et leur application aux quations intgrales}, Fund. Math., \textbf{3} (1922),133--181 (French).

\bibitem{BU} ] M. Bukatin, R. Kopperman, S. Matthews, and H. Pajoohesh, \emph{Partial Metric Spaces}, 
Amer. Math. Monthly, \textbf{116} (2009), 708--718.


\bibitem{CA}  H. Cakalli, A. S$\ddot{o}$nmez and C.Genc, \emph{On Equivalence of Topological Vector Space Valued Cone Metric Spaces and Metric spaces}, Appl. Math. Lett., \textbf{25} (2012), 429--433.

%\bibitem{CH} S.K. Chatterjee, \emph{Fixed point theorems}, Rend. Acad. Bulgare Sc., \textbf{25} (1972), 727--730.

\bibitem{CI}  L.B. $\acute{C}$iri$\acute{c}$, \emph{On some maps with a nonunique fixed point}, Publ. Inst. Math., \textbf{17} (1974), 52–-58. 

\bibitem{DAH} R.D. Daheriya, R. Jain and M. Ughade, \emph{Some unique fixed point theorems in partial cone metric spaces}, Inter. J. Fuzzy Math. Systems, \textbf{1}(2) (2011), 211--218.

%\bibitem{DE} D. Dey and M. Saha, \emph{Partial cone metric space and some fixed point theorems}, TWMS J. App. Eng. Math., \textbf{3}(1) (2013), 1--9.

\bibitem{HU}  L.-G. Huang, X. Zhang, \emph{Cone metric spaces and ﬁxed point theorems of contractive mappings}, J. Math. Anal. Appl., \textbf{332}(2) (2007), 1468--1476.

%\bibitem{KA} R. Kannan, \emph{Some results on ﬁxed points}, Bull. Calcutta Math. Soc., \textbf{60} (1968), 71--76.

\bibitem{KAR} E. Karapinar and I.M. Erhan, \emph{Fixed point theorems for operators on partial metric spaces}, Appl. Math. Lett., \textbf{24} (2011), 1894--1899.

\bibitem{KAR1} E. Karapinar, \emph{A new non-unique fixed point theorem}, Ann. Funct. Annal., \textbf{2}(1) (2011), 51--58.

\bibitem{LI} S. D. Lin, \emph{A common ﬁxed point theorem in abstract spaces}, Indian J. Pure Appl. Math., \textbf{18}(8) (1987), 685--690.

\bibitem{MAH}  S. K. Mahlotra, S. Shukla, R. Sen, N. Verma, \emph{Fixed point theorems in partial cone metric spaces}, Inter. J. Math. Arch., \textbf{2}(4)  (2011), 610--616.

\bibitem{MAT}  S. G. Matthews, \emph{Partial Metric Topology}, in: Proceedings of the 8th Summer Conference on Topology and its Applications, 728, Annals of The New york Academy of Sciences, (1994), 183--197. MR 98d:54054.

%\bibitem{RE} S. Reich, \emph{Kannan’s ﬁxed point thorem}, Boll. Un. Math. Ital., \textbf{4} (1971), 1--11.

\bibitem{RZ}  B. Rzepecki, \emph{On ﬁxed point theorems of Maia type}, Publ. Inst. Math., \textbf{28}(42) (1980), 179--186.

\bibitem{SO} A. S$\ddot{o}$nmez, \emph{Fixed point theorems in partial cone metric spaces}, arXiv:1101.2741v1 [math.GN]. 

\bibitem{TU} D. Turkoglu, M. Abuloha, \emph{Cone metric spaces and ﬁxed point theorems in diametrically contractive mappings}, Acta Math. Sin. Engl. Series, \textbf{26}(3) (2010), 489--496.

\bibitem{XI} H. L-Guang and Z. Xian, \emph{Cone metric spaces and fixed point theorems of contractive mappings}, J. Math. Anal. Appl., \textbf{332} (2007), 1468--1476.
\end{thebibliography}

\documentclass[12pt, reqno]{amsart}
\usepackage{amsmath, amsthm, amscd, amsfonts, amssymb, graphicx, color}
\usepackage[bookmarksnumbered, colorlinks, plainpages]{hyperref}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{corollary}[theorem]{Corollary}
\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}
\newtheorem{example}[theorem]{Example}

\theoremstyle{remark}
\newtheorem{remark}[theorem]{Remark}
\numberwithin{equation}{section}

\begin{document}

\title[Common fixed point results in partial cone metric spaces]
{Common fixed point results in partial cone metric spaces}

\author[Tayebeh Laal Shateri]{Tayebe Laal Shateri}
\address{Tayebe Laal Shateri \\ Department of Mathematics and Computer
Sciences, Hakim Sabzevari University, Sabzevar, P.O. Box 397, IRAN}
\email{ \rm t.shateri@hsu.ac.ir; shateri@ualberta.ca}
\thanks{*The corresponding author:
t.shateri@hsu.ac.ir ; shateri@ualberta.ca (Tayebe Laal Shateri)}
 \subjclass[2010] {Primary 47H10;
Secondary 54H25} \keywords{Fixed point, cone metric space, partial cone metric space, contraction.}
 \maketitle

\begin{abstract}
The main aim of this paper is to study of fixed point theory in partial cone metric spaces. Infact, some common fixed point theorems for two mappings in partial cone metric spaces are obtained.
 \vskip 3mm
\end{abstract}

\section{Introduction and preliminaries}\vskip 2mm
The concept of a partial metric, and any concept related to a partial metric play a very important role not only in pure mathematics but also in other branches of science involving mathematics especially in computer science, information science, and biological science.
In partial metric spaces, the self-distance for any point need not be equal to zero.\\
In 1994, Matthews \cite{MAT} introduced the notion of partial metric space. In 2009, Bukatin established the precise relationship between partial metric spaces and the so called weightable quasi-metric space \cite{BU}.  
In 1980, Rzepecki \cite{RZ} introduced a generalized metric $d_E$ on a set $\mathcal X$ in a way that $d_E:\mathcal X\times \mathcal X\to P$, replacing the set of real numbers with a Banach space $E$ where $P$ is a normal cone in $E$ with a partial order $\leqslant$. Many authors to investigate topological properties of cone metric spaces (see \cite{CA,LI,TU}).

Fixed point theorems are the basic mathematical tools used in showing the
existence of solution concept in such diverse ﬁelds as biology, chemistry, economics, engineering, and game theory. It is well-known that the Banach contraction principle \cite{BA} is a fundamental result in ﬁxed point theory, which has been used, and extended in many diﬀerent directions. Mattews proved the Banach ﬁxed point theorem in Partial metric space \cite{MAT}. Some fixed point theorems of contractive mappings for partial metric spaces and cone metric spaces have proved by many authors (see \cite{AB1,AB2,HU}). Partial cone metric space and fixed point theorem have investigated by Mahlotra, et al. \cite{MAH}.\\

In this paper, by using some ideas of \cite{DAH,KAR} we prove some common fixed point theorems for two mappings in partial cone metric spaces.

The following definitions and results will be needed in the sequel.
\begin{definition}(Partial metric space) A partial metric on a non-empty set $\mathcal X$ is a function $p:\mathcal X\times \mathcal X \to \mathbb R^+$ such that for all $x,y,z\in \mathcal X$ the followings hold\\
$(p_1)\; 0\leq p(x,x)\leq p(x,y),$\\
$(p_2)\; x=y \quad \text{if and only if}\quad p(x,x)=p(x,y)=p(y,y),$\\
$(p_3)\; p(x,y)=p(y,x),$\\
$(p_4)\; p(x,y)\leq p(x,z)+p(z,y)-p(z,z).$\\
Then the pair $(\mathcal X,p)$ is called a partial metric space. 
\end{definition}
It is clear that if $p(x,y)=0$, then $(p1)$ and $(p2)$ imply that $x=y$. But if $x=y$, $p(x,y)$ may not be $0$. A basic example of a partial metric space is the pair $(\mathbb R^+,p)$, where $p(x,y)=\max{x,y}$ for all $x,y\in \mathbb R^+$.\\
Let $E$ be a real Banach space and $P$ a subset of $E$. $P$ is called a cone if it satisfies the followings\\
$(C_1)\; P \;\text{is closed, non-empty and}\; P\neq \{0\},$\\
$(C_2)\; ax+by\in P \quad \text{for all}\; x,y\in P\; \text{and non-negative real numbers}\; a,b,$\\
$(C_3)\; P\cap (-P)=\{0\}.$\\
For a given cone $P\subseteq E$, we can deﬁne a partial ordering $\leqslant $ on $E$ with respect to $P$ by $x\leqslant y$ if and only if $y-x\in P$. We write $x<y$ to 
indicate that $x\leqslant y$ but $x\neq y$, while $x\ll y$ will stand for $y-x\in  intP$, in which $intP$ denotes the interior of $P$. The cone $P$ is called normal if there is a constant number $M>0$ such that for all $x,y\in E$ where $0\leqslant x\leqslant y$ implies $\|x\|\leqslant M\|y\|$. The least positive number satisfying above is called the normal constant of $P$. 

Let $E$ be a Banach space, $P$ a cone in $E$ with $intP\neq\Phi$ and $\leqslant$ is partial ordering with respect to $P$. 
\begin{definition}(Cone metric space) Let $\mathcal X$ be a non-empty set. The  mapping \\$d:\mathcal X\times\mathcal X\to E$ is said to be a cone metric on $\mathcal X$ if for all $x,y,z\in X$ the followings hold\\
$(CM_1)\; 0\leqslant d(x,y)\;\text{and}\; d(x,y)=0\; \text{if and only if}\; x=y,$\\
$(CM_2)\; d(x,y)=d(y,x),$\\
$(CM_3)\; d(x,y)\leqslant d(x,z)+d(z,y).$\\
Then $(\mathcal X,d)$ is called a cone metric space.
\end{definition} 
Mahlotra and et al. \cite{MAH} and S$\ddot{o}$nmez \cite{SO} introduced the notion of partial cone metric space and its topological characterization. We now state the deﬁnition of partial cone metric space.
\begin{definition} (Partial cone metric space) A partial cone metric on a non-empty set $\mathcal X$ is a function $p:\mathcal X\times\mathcal X\to E$ such that for all $x,y,z\in X$\\ 
$(PCM_1)\; 0\leqslant p(x,x)\leqslant p(x,y),$\\
$(PCM_2)\; x=y\quad \text{if and only if}\quad p(x,x)=p(x,y)=p(y,y),$\\
$(PCM_3)\; p(x,y)=p(y,x),$\\
$(PCM_4)\; p(x,y)\leqslant p(x,z)+p(z,y)-p(z,z).$
\end{definition}
A partial cone metric space is a pair $(\mathcal X,p)$ such that $\mathcal X$ is a non-empty set and $p$ is a partial cone metric on $\mathcal X$. It is clear that, if $p(x,y)=0$, then $(PCM_1)$ and $(PCM_2)$ imply that $x=y$. But the converse is not true ingeneral. A cone metric space is a partial cone metric space, but there exists partial cone metric spaces which are not cone metric space. we give the following example from \cite{SO}.
\begin{example}
Let $E=\mathbb R^2$, $P=\{(x,y)\in E:\; x,y\geqslant 0\}$ and $\mathcal X=\mathbb R^+$ and $p:\mathcal X\times \mathcal X\to E$ deﬁned by $p(x,y)=(\max\{x,y\},k\max\{x,y\})$ where $k\geqslant 0$ is a constant. Then $(\mathcal X,p)$ is a partial cone metric space which is not a cone metric space.
\end{example}
\begin{remark}\label{1}
Suppose $(\mathcal X,p)$ is a partial cone metric space, then
$$d(x,y)=2p(x,y)-p(x,x)-p(y,y)\;,\quad  {\text for\; all}\; x,y\in \mathcal X$$
deﬁnes a cone metric on X.
\end{remark}
Following, We give some properties of partial cone metric spaces, for more details see \cite{SO}.
\begin{theorem}
Every partial cone metric space $(\mathcal X,p)$ is a topological space.
\end{theorem}
\begin{definition}
Let $(\mathcal X,p)$ be a partial cone metric space. Let $\{x_n\}$ be a sequence in $\mathcal X$ and $x\in \mathcal X$.\\
$(i)$ $\{x_n\}$ is said to be convergent to $x$ and $x$ is called a limit of $\{x_n\}$ if
$$\lim_{n\to\infty}p(x_n,x)=\lim_{n\to\infty}p(x_n,x_n)=p(x,x).$$
$(ii)$ Let $(\mathcal X,p)$ be a partial cone metric space. $\{x_n\}$ be a sequence in$\mathcal  X$. $\{x_n\}$ is Cauchy sequence if there is $x\in P$ such that for every $\varepsilon >0$ there is $N$ such that for all $n,m> N$, $\|p(x_n,x_m)-x\|<\varepsilon$.\\
$(i)\; (\mathcal X,p)$ is said to be complete if every Cauchy sequence in $(\mathcal X,p)$ is convergent in $(\mathcal X,p)$.
\end{definition}
%---------------------------------------------------------------------------------------------------------------------------------------------------------
\section{ Main results}
In this section we give some common fixed point results in partial cone metric spaces. We assume that $(\mathcal X,p)$ be a complete partial cone metric space, $P$ be a normal cone with normal constant $M$.
%------------------------------------------------------------------------------
\begin{theorem}\label{th1}
Let $T$ and $S$ be two mappings defined on $\mathcal X$ satisfy the following contraction
\begin{equation}\label{eq1.1}
p(Tx,Sy)\leqslant\alpha p(x,Tx)+\beta p(y,Sy),
\end{equation}
for all $x,y\in \mathcal X$, and $\alpha+\beta<1,\; \alpha,\beta\in[0,1)$. Then $T$ and $S$ have a unique common fixed point.
\begin{proof}
Let $x_0\in \mathcal X$ be arbitrary. Put $Tx_0=x_1$ and $Sx_1=x_2$, then \eqref{eq1.1} implies that
\begin{align*}
p(x_1,x_2)&=p(Tx_0,Sx_1)\leqslant \alpha p(x_0,Tx_0)+\beta p(x_1,Sx_1)\\
&=\alpha p(x_0,x_1)+\beta p(x_1,x_2),
\end{align*}
hence $p(x_1,x_2)\leqslant\frac{\alpha}{1-\beta}p(x_0,x_1)$, where $0<K=\frac{\alpha}{1-\beta}<1$. Continuing this process we obtain a sequence $\{x_n\}$ in $\mathcal X$ such that $x_{2n-1}=Tx_{2n-2},x_{2n}=Sx_{2n-1}$ and $$p(x_{n+1},x_n)\leqslant Kp(x_n,x_{n-1}),$$ which implies that $$p(x_{n+1},x_n)\leqslant K^np(x_1,x_0).$$
Now, we show that the sequence $\{x_n\}$ is Cauchy. For $m>n$ we have
\begin{align*}
p(x_m,x_n)&\leqslant p(x_m,x_{m-1})+p(x_{m-1},x_{m-2})\\&+\cdots+p(x_{n+1},x_n)-\sum_{i=1}^{m-n-1}p(x_{n-i},x_{n-i})\\
&\leqslant\big[K^{m-1}+K^{m-2}+\cdots +K^n\big]p(x_1,x_0)\\
&\leqslant K^n\frac{1-K^{m-n}}{1-K}p(x_1,x_0)\\
&\leqslant\frac{K^n}{1-K}p(x_1,x_0),
\end{align*}
hence $\|p(x_m,x_n)\|\leq M\frac{K^n}{1-K}\|p(x_1,x_0)\|\to0$ as $m,n\to\infty$. Thus $\{x_n\}$ is a Cauchy sequence such that $\lim_{m\to\infty}p(x_m,x_n)=0$. Since $(\mathcal X,p)$ is a complete partial cone metric space, so there exists $x\in \mathcal X$ such that $$p(x,x)=\lim_{n\to\infty}p(x_n,x)=\lim_{n\to\infty}p(x_n,x_n)=0.$$
From \eqref{eq1.1} we get
\begin{align*}
p(Tx,x)&\leqslant p(Tx,x_{2n})+p(x_{2n},x)-p(x_{2n},x_{2n})\\
&\leqslant p(Tx,Sx_{2n-1})+p(x_{2n},x)\\
&\leqslant\alpha p(x,Tx)+\beta p(x_{2n-1},Sx_{2n-1})+p(x_{2n},x)\\
&\leqslant\alpha p(x,Tx)+\beta p(x_{2n-1},x_{2n})+p(x_{2n},x),
\end{align*}
hence $p(Tx,x)\leqslant\frac{1}{1-\alpha}\big[\beta p(x_{2n-1},x_{2n})+p(x_{2n},x)\big]$ and so $$\|p(Tx,x)\|\leqslant\frac{M}{1-\alpha}\big[\beta\|p(x_{2n-1},x_{2n})+p(x_{2n},x)\|\big]\to 0.$$
Therefore $p(Tx,x)=0$, similarly, we get $p(Sx,x)=0$. 
Thus 
$$0\leq d_p(x,Tx)=2p(x,Tx)-p(x,x)-p(Tx,Tx)\leq 0,$$ and $$0\leq d_p(x,Sx)=2p(x,Sx)-p(x,x)-p(Sx,Sx)\leq 0,$$
hence $d_p(x,Tx)=d_p(x,Sx)=0$. Since $d_p$ is a metric, hence $Tx=x=Sx$, 
%$p(Tx,Sx)\leqslant\alpha p(x,Tx)+\beta p(x,Sx)=0$ and $p(Tx,Tx)=p(Sx,Sx)=0$. %Consequently
%$$p(Tx,Tx)=p(Tx,x)=p(x,x)=0\; \text{and}\quad p(Sx,Sx)=p(Sx,x)=p(x,x)=0.$$
and so $x$ is the common fixed point of $T,S$.\\
{\bf Uniquness}. Let $y$ be another common fixed point of $T,S$, then $p(x,x)=p(y,y)=0$ and
$$p(x,y)=p(Tx,Sy)\leqslant\alpha p(x,Tx)+\beta p(y,Sy)=\alpha p(x,x)+\beta p(y,y)=0,$$
and so $p(x,y)=0=p(x,x)=p(y,y)$, therefore $x=y$. This completes the proof.
\end{proof}
\end{theorem}
%------------------------------------------------------------------------------
If in Theorem \ref{th1} we set $0<\beta=\alpha<\frac{1}{2}$, then we get the following result.
\begin{corollary}\label{cor1}
Let $T$ and $S$ be two mappings defined on $\mathcal X$ satisfy the following contraction
\begin{equation}\label{eqcor1}
p(Tx,Sy)\leqslant\alpha\big[p(x,Tx)+p(y,Sy)\big],
\end{equation}
for all $x,y\in \mathcal X$. Then $T$ and $S$ have a unique common fixed point.
\end{corollary}
%---------------------------------------------------------------------
Using an idea of \cite{XI}, we give an example of mappings on a partial cone metric space which satisfy in the conditions of Corollary \ref{cor1}.
\begin{example}
Let $E=l^1$, and $P=\{\{x_n\}:x_n\geqslant 0\}$ a normal cone. 
%Let $(\mathbb R^+)^{\omega}$ be the set of all inﬁnite sequences over a set $%\mathbb R^+$, 
Set \\$\mathcal X=\{\{x_n\}:x_n\in [0,\frac{\pi}{4}], \sum_nx_n<\infty\}$ and $p:\mathcal X\times \mathcal X\to E$ deﬁned by
$$p(x,y)=(x_1 \vee y_1,x_2 \vee y_2,\ldots,x_n \vee y_n,\ldots)$$
where the symbol $\vee$ denotes the maximum, i.e. $x \vee y = \max\{x,y\}$. Then $(X,p)$ is a partial cone metric space, (which is not a cone metric space). Consider $T,S:\mathcal X\to \mathcal X$ defined by $$T(\{x_n\})=\{\frac{x_n\tan{x_n}}{3}\}\quad {\text and}\quad S(\{x_n\})=\{\frac{x_n}{4}\}.$$ Then 
\begin{align*}
p\Big(T(\{x_n\}),S(\{y_n\})\Big)&=p\Big(\{\frac{x_n\tan{x_n}}{3}\},\{\frac{y_n}{4}\}\Big)\\
&=\Big(\frac{x_1\tan{x_1}}{3}\vee \frac{y_1}{4},\frac{x_2\tan{x_2}}{3}\vee \frac{y_2}{4},\ldots,\frac{x_n\tan{x_n}}{3}\vee \frac{y_n}{4},\ldots\Big)\\
&\leq\frac{1}{3}\Big[p\Big(\{x_n\},T(\{x_n\})\Big)+p\Big(\{y_n\},S(\{y_n\})\Big)\Big]\\
&=\{\frac{x_n+y_n}{3}\}.
\end{align*}
Therefore $T$ and $S$ satisfy the contractive condition \eqref{eqcor1} with $\alpha=\frac{1}{3}$. The common fixed point of $T$ and $S$ is $0$.
\end{example}
%------------------------------------------------------------------------------
We present the following common fixed point theorem for two mappings satisfying contractive condition in partial cone metric spaces.
\begin{theorem}\label{th2}
Let $T$ and $S$ be two mappings defined on $\mathcal X$ satisfy the following contraction
\begin{equation}\label{eq2.1}
p(Tx,Sy)\leqslant\alpha p(x,y)+\beta p(x,Tx)+\gamma p(y,Sy),
\end{equation}
for all $x,y\in \mathcal X$, and $\alpha+\beta+\gamma<1,\; \alpha,\beta,\gamma\in[0,1)$. Then $T$ and $S$ have a unique common fixed point.
\begin{proof}
Let $\{x_n\}$ be as sequence in the proof of Theorem \ref{th1}, then \eqref{eq2.1} implies that
\begin{align*}
p(x_1,x_2)&=p(Tx_0,Sx_1)\leqslant \alpha p(x_0,x_1)+\beta p(x_0,Tx_0)+\gamma p(x_1,Sx_1)\\
&=\alpha p(x_0,x_1)+\beta p(x_0,x_1)+\gamma p(x_1,x_2),
\end{align*}
hence $p(x_1,x_2)\leqslant\frac{\alpha+\beta}{1-\gamma}p(x_0,x_1)$, where $0<K=\frac{\alpha+\beta}{1-\beta}<1$. Continuing this process we obtain a sequence $\{x_n\}$ in $\mathcal X$ such that $x_{2n-1}=Tx_{2n-2},x_{2n}=Sx_{2n-1}$ and $$p(x_{n+1},x_n)\leqslant K^np(x_1,x_0).$$
As the proof of Theorem \ref{th1}, one can show that the sequence $\{x_n\}$ is Cauchy in $(\mathcal X,p)$. Since $(\mathcal X,p)$ is a complete partial cone metric space, so there exists $x\in \mathcal X$ such that $$p(x,x)=\lim_{n\to\infty}p(x_n,x)=\lim_{n\to\infty}p(x_n,x_n)=0.$$
From \eqref{eq2.1} we get
\begin{align*}
p(Tx,x)&\leqslant p(Tx,x_{2n})+p(x_{2n},x)-p(x_{2n},x_{2n})\\
&\leqslant p(Tx,Sx_{2n-1})+p(x_{2n},x)\\
&\leqslant\alpha p(x,x_{2n-1})+\beta p(x,Tx)+\gamma p(x_{2n-1},Sx_{2n-1})+p(x_{2n},x)\\
&\leqslant\alpha p(x,x_{2n-1})+\beta p(x,Tx)+\gamma p(x_{2n-1},x_{2n})+p(x_{2n},x),
\end{align*}
hence $p(Tx,x)\leqslant\frac{1}{1-\beta}\big[\alpha p(x,x_{2n-1})+\gamma p(x_{2n-1},x_{2n})+p(x_{2n},x)\big]$ and so $$\|p(Tx,x)\|\leqslant\frac{M}{1-\beta}\big[\alpha\|p(x,x_{2n-1})\|+\gamma\|p(x_{2n-1},x_{2n})\|+\|p(x_{2n},x)\|\big]\to 0.$$
Therefore $p(Tx,x)=0$, Similarly, $p(Sx,x)=0$. 
%and 
%$$p(Tx,Sx)\leqslant\alpha p(x,x)+\beta p(x,Tx)+\gamma p(x,Sx)=0$$ also  %$p(Tx,Tx)=p(Sx,Sx)=0$. Consequently
%$$p(Tx,Tx)=p(Tx,x)=p(x,x)=0\; \text{and}\quad p(Sx,Sx)=p(Sx,x)=p(x,x)=0.$$
%$$0\leq d_p(x,Tx)=2p(x,Tx)-p(x,x)-p(Tx,Tx)\leq 0,$$ and $$0\leq d_p(x,Sx)=2p(x,Sx)-p(x,x)-p(Sx,Sx)\leq 0,$$
As the proof of Theorem \ref{th1}, we have $d_p(x,Tx)=d_p(x,Sx)=0$, hence $Tx=x=Sx$, and so $x$ is the common fixed point of $T,S$. As the proof of Theorem \ref{th1}, the fixed point is unique and the proof is completed.
\end{proof}
\end{theorem}
%------------------------------------------------------------------------------
If in Theorem \ref{th2} we set $0<\alpha=\beta=\gamma<\frac{1}{3}$, then we get the following result.
\begin{corollary}
Let $T$ and $S$ be two mappings defined on $\mathcal X$ satisfy the following contraction
\begin{equation}
p(Tx,Sy)\leqslant\alpha\big[p(x,y)+p(x,Tx)+p(y,Sy)\big],
\end{equation}
for all $x,y\in \mathcal X$. Then $T$ and $S$ have a unique common fixed point.
\end{corollary}
%------------------------------------------------------------------------------
Next, we present the following common fixed point theorem for two mappings defined on partial cone metric spaces with contractive condition of rational type. 
\begin{theorem}\label{th3}
Let $T$ and $S$ be two mappings defined on $\mathcal X$ satisfy the following contraction
\begin{equation}\label{eq3.1}
p(Tx,Sy)\leqslant\alpha\frac{p(x,Tx)p(y,Sy)}{1+p(x,y)}+\beta p(x,y),
\end{equation}
for all $x,y\in \mathcal X$, and $\alpha+\beta<1,\; \alpha,\beta\in[0,1)$. Then $T$ and $S$ have a unique common fixed point.
\begin{proof}
Let $\{x_n\}$ be as sequence in the proof of Theorem \ref{th1}. Then \eqref{eq3.1} implies that
\begin{align*}
p(x_1,x_2)&=p(Tx_0,Sx_1)\\&\leqslant \alpha\frac{p(x_0,Tx_0)p(x_1,Sx_1)}{1+p(x_0,x_1)}+\beta p(x_0,x_1)\\
&=\alpha\frac{p(x_0,x_1)p(x_1,x_2)}{1+p(x_0,x_1)}+\beta p(x_0,x_1)\\
&<\alpha p(x_1,x_2)+\beta p(x_0,x_1),
\end{align*}
hence $p(x_1,x_2)\leqslant\frac{\beta}{1-\alpha}p(x_0,x_1)$, where $0<K=\frac{\beta}{1-\alpha}<1$. Continuing this process we obtain a sequence $\{x_n\}$ in $\mathcal X$ such that $x_{2n-1}=Tx_{2n-2},x_{2n}=Sx_{2n-1}$ and $$p(x_{n+1},x_n)\leqslant K^np(x_1,x_0).$$
As the proof of Theorem \ref{th1}, it is shown that the sequence $\{x_n\}$ is Cauchy, so there exists $x\in \mathcal X$ such that $$p(x,x)=\lim_{n\to\infty}p(x_n,x)=\lim_{n\to\infty}p(x_n,x_n)=0.$$
From \eqref{eq3.1} we get
\begin{align*}
p(Tx,x)&\leqslant p(Tx,x_{2n})+p(x_{2n},x)-p(x_{2n},x_{2n})\\
&\leqslant p(Tx,Sx_{2n-1})+p(x_{2n},x)\\
&\leqslant\alpha\frac{p(x,Tx)p(x_{2n-1},Sx_{2n-1})}{1+p(x,x_{2n-1})}+\beta p(x,x_{2n-1})+p(x_{2n},x)\\
&\leqslant\alpha\frac{p(x,Tx)p(x_{2n-1},x_{2n})}{1+p(x,x_{2n-1})}+\beta p(x,x_{2n-1})+p(x_{2n},x),
\end{align*}
hence $$\|p(Tx,x)\|\leqslant M\frac{1+p(x,x_{2n-1})}{\|1-\alpha p(x_{2n-1},x_{2n})\|}\big[\beta\|p(x,x_{2n-1})\|+\|p(x_{2n},x)\|\big]\to 0.$$
Therefore $p(Tx,x)=0$. Similarly, $p(Sx,x)=0$.  
%also 
%$$p(Tx,Sx)\leqslant\alpha\frac{p(x,Tx)p(x,Sx)}{1+p(x,x)}+\beta p(x,x)=0$$ and $p(Tx,Tx)=p(Sx,Sx)=0$. Thus $$p(Tx,Tx)=p(Tx,x)=p(x,x)=0\; \text{and}\quad p(Sx,Sx)=p(Sx,x)=p(x,x)=0.$$
%Thus
%$$0\leq d_p(x,Tx)=2p(x,Tx)-p(x,x)-p(Tx,Tx)\leq 0,$$ and $$0\leq %d_p(x,Sx)=2p(x,Sx)-p(x,x)-p(Sx,Sx)\leq 0,$$
As the proof of Theorem \ref{th1} we deduce $d_p(x,Tx)=d_p(x,Sx)=0$. So $Tx=x=Sx$, that is  $x$ is the common fixed point of $T,S$. It is easy to see that the common fixed point is unique and so the proof is completed.
\end{proof}
\end{theorem}
%------------------------------------------------------------------------------
Now, we give an example of mappings on a partial cone metric space which satisfy in the conditions of Theorem \ref{th3}.
\begin{example}\label{ex1}
Let $E=\mathbb R^2$, $P=\{(x,y)\in E:x,y\geqslant 0\}$, $\mathcal X=[0,\frac{\pi}{4}]$, and $p:\mathcal X\times \mathcal X\to E$ defined by $p(x,y)=(\max\{x,y\}, k\max\{x,y\})$ where $k>0$ is a constant. Define $$T,S:\mathcal X\to\mathcal X \quad{\text as}\quad T(x)=\frac{x}{2}\quad {\text and}\quad S(x)=\frac{x\sin x}{3}.$$ 
Let $x\leq y$, 
then 
$$p(x,y)=(y,ky),p(x,Tx)=(x,kx),p(y,Sy)=(y,ky).$$ Hence with $\frac{1}{2}\leq\alpha+\beta<1$, we get
\begin{align*}
\alpha\frac{p(x,Tx)p(y,Sy)}{1+p(x,y)}+\beta p(x,y)&=
\alpha\frac{(x,kx)(y,ky)}{1+(y,ky)}+\beta(y,ky)\\&\geqslant\beta(y,ky)\\
&\geqslant (\max\{\frac{x}{2},\frac{y\sin y}{3}\}, k\max\{\frac{x}{2},\frac{y\sin y}{3}\})\\&=p(Tx,Sy).
\end{align*}
Therefore $T$ and $S$ satisfy the conditions of Theorem \ref{th3}. The common fixed point of $T,S$ is $x=0$.
\end{example}
%------------------------------------------------------------------------------
The next result has an analog on metric spaces and partial metric spaces, (see \cite{KAR,KAR1}).
\begin{theorem}\label{th4}
Let $T$ and $S$ be two mappings defined on $\mathcal X$ satisfy the following contraction
\begin{equation}\label{eq4.1}
\alpha p(Tx,Sy)+\beta\big[p(x,Tx)+p(y,Sy)\big]+\gamma\big[p(Tx,y)+p(x,Sy)\big]\leqslant sp(x,y)+rp(x,STx),
\end{equation}
for all $x,y\in \mathcal X$, and the constants $\alpha,\beta,\gamma,s,r$, where 
$$0\leq \frac{s-\beta}{\alpha+\beta}<1,\; \alpha+\beta\neq 0,\; \alpha+\beta+\gamma>0,\; \gamma>0,\;{\text and}\quad \gamma-r\geqslant 0.$$
Then $T$ and $S$ have a common fixed point.
\begin{proof}
Let $\{x_n\}$ be as sequence in the proof of Theorem \ref{th1}, we deduce from \eqref{eq4.1} that
\begin{align*}
\alpha p(Tx_0,Sx_1)&+\beta\big[p(x_0,Tx_0)+p(x_1,Sx_1)\big]+\gamma\big[p(Tx_0,x_1)+p(x_0,Sx_1)\big]\\&\leqslant sp(x_0,x_1)+rp(x_0,STx_0),
\end{align*}
hence 
\begin{align*}
\alpha p(x_1,x_2)&+\beta\big[p(x_0,x_1)+p(x_1,x_2)\big]+\gamma\big[p(x_1,x_1)+p(x_0,x_2)\big]\\&\leqslant sp(x_0,x_1)+rp(x_0,Sx_1),
\end{align*}
so $$(\alpha+\beta)p(x_1,x_2)+(\gamma-r)p(x_0,x_2)+\gamma p(x_1,x_1)\leqslant(s-\beta)p(x_0,x_1).$$
using the fact that $(\gamma-r)p(x_0,x_2)+\gamma p(x_1,x_1)\geqslant 0$, we get
$$p(x_1,x_2)\leqslant\frac{s-\beta}{\alpha+\beta}p(x_0,x_1)=Kp(x_0,x_1),$$
where $0\leq K=\frac{s-\beta}{\alpha+\beta}<1$. Continuing this process we have $$p(x_n,x_{n+1})\leq K^n p(x_0,x_1).$$
As the proof of Theorem \ref{th1}, it is shown that the sequence $\{x_n\}$ is Cauchy, so there exists $x\in \mathcal X$ such that 
\begin{equation}\label{eq4.2}
p(x,x)=\lim_{n\to\infty}p(x_n,x)=\lim_{n\to\infty}p(x_n,x_n)=0.
\end{equation}
From \eqref{eq4.1} we get
\begin{align*}
\alpha p(Tx_{2n},Sx)&+\beta\big[p(x_{2n},Tx_{2n})+p(x,Sx)\big]+\gamma\big[p(Tx_{2n},x)+p(x_{2n},Sx)\big]\\&\leqslant sp(x_{2n},x)+rp(x_{2n},STx_{2n}),
\end{align*}
hence 
\begin{align*}
\alpha p(x_{2n+1},Sx)&+\beta\big[p(x_{2n},x_{2n+1})+p(x,Sx)\big]+\gamma\big[p(x_{2n+1},x)+p(x_{2n},Sx)\big]\\&\leqslant sp(x_{2n},x)+rp(x_{2n},x_{2n+2}),
\end{align*}
thus
\begin{align*}
p(x,Sx)&\leqslant \frac{1}{\beta}\Big[sp(x_{2n},x)+rp(x_{2n},x_{2n+2})-\alpha p(x_{2n+1},Sx)\\&-\beta p(x_{2n},x_{2n+1})-\gamma\big[p(x_{2n+1},x)+p(x_{2n},Sx)\big]\Big]\\
&\leqslant \frac{1}{\beta}\Big[sp(x_{2n},x)+rp(x_{2n},x_{2n+2})-\beta p(x_{2n},x_{2n+1})-\gamma p(x_{2n+1},x)\Big]
\end{align*}
therefore
\begin{align*}
\|p(x,Sx)\|&\leqslant \frac{M}{\beta}\|sp(x_{2n},x)+rp(x_{2n},x_{2n+2})\beta p(x_{2n},x_{2n+1})-\gamma p(x_{2n+1},x)\|
\end{align*}
and using \eqref{eq4.2} we obtain $p(x,Sx)=0$, similarly we get $p(x,Tx)=0$. 
%Thus 
%$$0\leq d_p(x,Tx)=2p(x,Tx)-p(x,x)-p(Tx,Tx)\leq 0,\quad 0\leq d_p(x,Sx)=2p(x,Sx)-p(x,x)-p(Sx,Sx)\leq 0,$$
As the proof of Theorem \ref{th1}, we get $d_p(x,Tx)=d_p(x,Sx)=0$. Thus $Tx=x=Sx$, which completes the proof. 
\end{proof}
\end{theorem}
%------------------------------------------------------------------------------
\begin{theorem}\label{th5}
Let $T$ and $S$ be two mappings defined on $\mathcal X$ satisfy the following contraction
\begin{equation}\label{eq5.1}
p(Tx,Sy)\leqslant\alpha\max\{p(x,y),p(x,Tx),p(y,Sy)\}
\end{equation}
for all $x,y\in \mathcal X$, and $\alpha\in [0,1)$. Then $T$ and $S$ have a unique common fixed point.
\begin{proof}
Let $\{x_n\}$ be as sequence in the proof of Theorem \ref{th1}, we deduce from \eqref{eq5.1} that
\begin{align*}
p(Tx_{2n-2},Sx_{2n-1})\leqslant\alpha\max\{p(x_{2n-2},x_{2n-1}),p(x_{2n-2},Tx_{2n-2}),p(x_{2n-1},Sx_{2n-1})\},
\end{align*}
hence 
\begin{align*}
p(x_{2n-1},x_{2n})&\leqslant\alpha\max\{p(x_{2n-2},x_{2n-1}),p(x_{2n-2},x_{2n-1}),p(x_{2n-1},x_{2n})\}\\
&=\alpha\max\{p(x_{2n-2},x_{2n-1}),p(x_{2n-1},x_{2n})\}\\
&\leqslant\alpha p(x_{2n-2},x_{2n-1})\\
&\leqslant\alpha^n p(x_0,x_1).
\end{align*}
Now, for $m>n$ we have $$p(x_m,x_n)\leqslant\frac{\alpha^n}{1-\alpha}p(x_0,x_1).$$
hence $\|p(x_m,x_n)\|\leqslant M\frac{\alpha^n}{1-\alpha}\|p(x_0,x_1)\|$, thus the sequence $\{x_n\}$ is Cauchy in $(\mathcal X,p)$ and so there exists $x\in \mathcal X$ such that
\begin{equation}\label{eq5.2}
p(x,x)=\lim_{n\to\infty}p(x_n,x)=\lim_{n\to\infty}p(x_n,x_n)=0.
\end{equation}
From \eqref{eq5.1} we get
\begin{align*}
p(Tx,x)&\leqslant p(Tx,Sx_{2n-1})+p(Sx_{2n-1},x)-p(Sx_{2n-1},Sx_{2n-1})\\
&\leqslant\alpha\max\{p(x,x_{2n-1}),p(x,Tx),p(x_{2n-1},x_{2n})\}+p(x_{2n},x)\\
&\leqslant\alpha\max\{p(x,x_{2n-1}),p(x_{2n-1},x_{2n})\}+p(x_{2n},x),
\end{align*}
therefore $\|p(Tx,x)\|\leqslant M\alpha\|\max\{p(x,x_{2n-1}),p(x_{2n-1},x_{2n})\}+p(x_{2n},x)\|\to 0$. So 
$p(x,Tx)=0$. Similarly we deduce that $p(x,Sx)=0$. Also \eqref{eq5.1} implies that
\begin{align*}
p(Tx,Sx)\leqslant\alpha\max\{p(x,x),p(x,Tx),p(x,Sx)\}=0,
\end{align*}
we have that $p(x,x)=p(x,Tx)=p(x,Sx)=p(Tx,Sx)=0$, which implies that $Tx=Sx=x$ and so $x$ is the common fixed point of $T$ and $S$. Now if $y$ is another common ﬁxed point of $T,S$, then 
$$p(x,y)=p(Tx,Sy)\leqslant\alpha\max\{p(x,y),p(x,Tx),p(y,Sy)\}=0,$$ we get $p(x,y)=p(x,x)=p(y,y)=0$, this completes the proof. 
\end{proof}
\end{theorem}
%-----------------------------------------------------------------------------
Final, we give an example which satisfy in the conditions of Theorem \ref{th5}.
\begin{example}
Let $E,P$ and $\mathcal X=[0,\frac{\pi}{4}]$. Define $p:\mathcal X\times \mathcal X\to E$ by $p(x,y)=(\max\{x,y\}, k\max\{x,y\})$ where $k\geqslant 0$ is a constant. Define $$T,S:\mathcal X\to\mathcal X \quad{\text as}\quad T(x)=\frac{x(1-\cos x)}{3}\quad {\text and}\quad S(x)=\frac{x}{2}.$$ 
Let $x\leq y$, then 
$$p(x,y)=(y,ky),p(x,Tx)=(x,kx),p(y,Sy)=(y,ky),p(Tx,Sy)=(\frac{y}{2},k\frac{y}{2}).$$ 
Hence $$\max\{p(x,y),p(x,Tx),p(y,Sy)\}=\max\{(y,ky),(x,kx),(y,ky)\}=(y,ky),$$
and so
\begin{equation}\label{ex1}
p(Tx,Sy)=\max\{\frac{x(1-\cos x)}{3},\frac{y}{2}\}=(\frac{y}{2},k\frac{y}{2})\leqslant\alpha(y,ky).
\end{equation}
If $x\geq y$, then
$$p(x,y)=(x,kx),p(x,Tx)=(x,kx),p(y,Sy)=(y,ky),$$
hence $$\max\{p(x,y),p(x,Tx),p(y,Sy)\}=\max\{(y,ky),(x,kx),(y,ky)\}=(x,kx),$$
thus
\begin{equation}\label{ex2}
p(Tx,Sy)=\max\{\frac{x(1-\cos x)}{3},\frac{y}{2}\}\leqslant\alpha(x,kx).
\end{equation}
\ref{ex1} and \ref{ex2} imply that $T$ and $S$ satisfy the conditions of Theorem \ref{th5} with $\alpha=\frac{2}{3}$. The common fixed point of $T,S$ is $x=0$.
\end{example}
%-----------------------------------------------------------------------------

\end{document}